# Tensor products of bimodules and bimodule pairs over monoids

### v2

@@ 5.1. Defining $f_{1}\otimes f_{2}$ via simple tensors
Here is the part that we can salvage in general.
@@
Suppose $\mathcal{M}_{1},\mathcal{N}_{1}$ are $\mathcal{T}$-modules and $\mathcal{M}_{2},\mathcal{N}_{2}$ are left $\mathcal{T}$-modules. If $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ is a right multiplicative map and $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$ is a left multiplicative map, then there is a 1-balanced map denoted as $f_{1}\times_{\circ}f_{2}:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes\mathcal{N}_{2}$ given by $(b_{1},b_{2})\mapsto f_{1}(b_{1})\otimes f_{2}(b_{2}).$
@@
Given a $(\mathcal{T}_{1},\mathcal{T})$-bimodule prepair $({\mathcal{M}_{1}},{\mathcal{M}_{1}}_{0})$, a $\mathcal{T},\mathcal{T}_{2}$-bimodule prepair $({\mathcal{N}},{\mathcal{N}}_{0})$, and a $\mathcal{T}_{1},\mathcal{T}_{2}$-bimodule prepair $(\mathcal{N},\mathcal{N}_{0})$, define $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{N})$ to be the set of $\mathcal{T}_{1},\mathcal{T}_{2}$ multiplicative maps $f:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}$ (as in Remark 5.6) satisfying the property that if $b_{j},b_{i,j}\in\mathcal{M}_{i}$ with $\sum_{j}b_{i,j}\in{\mathcal{M}_{i}}_{0},$ then
@@ 5.1.1. The tensor product of a weak morphism and a homomorphism when $\mathcal{M}_{2}$ is free
Proposition 5.5 raises the issue of how to define $(f\otimes g)(v\otimes w)$, with the hope of extending it to all of $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$.
When $\mathcal{M}_{2}$ is free as a $\mathcal{T}$-module, we can still define the tensor product of a weak morphism $f_{1}$ and a homomorphism $f_{2}$. Then we can write tensors uniquely in terms of a base, by taking an arbitrary element $\sum_{i\in I}v_{i}\otimes b_{i}$ as in Corollary 4.17(v), and put $(f\otimes g)(\sum_{i\in I}v_{i}\otimes b_{i})=f(v_{i})\otimes g(b_{i})$. We can phrase this as:
Proposition 5.8.
Notation as in Proposition 4.13, suppose $\mathcal{M}_{2}$ is free, with base $\{b_{i}:i\in I\}$. Then for any weak $(\mathcal{T}_{1},\mathcal{T})$-bimodule morphism (resp.  $\preceq$-morphism) $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ and any $(\mathcal{T},\mathcal{T}_{2})$-bimodule homomorphism $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$, there is a weak $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule morphism (resp.  $\preceq$-morphism)
@@
We do it for $\preceq$-morphisms. The proof for weak morphisms is parallel. Define the map $f_{1}\times_{\circ}f_{2}$ as in Proposition 5.5. Then $f_{1}\otimes f_{2}$ is well-defined when we write vectors of $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ uniquely in the form $\sum v_{i,j}\otimes b_{j}$, and
@@
Remark 5.9.
@@ 5.1.2. The tensor product of two weak morphisms or two $\preceq$-morphisms when $\mathcal{M}_{2}$ is free
In this case we can also modify the idea of [20], introducing the empty set and partial maps.
@@ 5.2. $\preceq$-Tensor products of $\preceq$-morphisms
We need to show that $f(\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2})\preceq f(\sum v_{i,1}\otimes w_{i,1})+f(\sum v_{i,2}\otimes w_{i,2})$. But if $\sum v_{i,1}^{\prime}\otimes w_{i^{\prime},1}=\sum v_{i,1}\otimes w_{i,1}$ and $\sum v_{i^{\prime},2}\otimes w_{i^{\prime},2}=\sum v_{i,2}\otimes w_{i,2}$ then
$\sum v_{i,1}^{\prime}\otimes w_{i^{\prime},1}+\sum v_{i^{\prime},2}\otimes w_{i^{\prime},2}=\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2},$
@@
Theorem 5.11 can be applied to hypermagmas in view of Remark 3.5, but at cost of introducing the empty set (the $\infty$ element of $\mathcal{P}(\mathcal{H}),$ cf. footnote 3 of Definition 3.1). In other words, $f$ could be undefined on tensors which can be rewritten in inequivalent ways.
@@ 5.2.1. Digression: Hypertensor products of maps
An intriguing (but problematic) way to obtain a theory that runs smoothly is by introducing a hyper-structure.
@@
Given maps $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$, $i=1,2,$ one could try defining $f:=f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{P}(\mathcal{N}_{1}\otimes\mathcal{N}_{2})$ by $f(\sum v_{i}\otimes w_{i})=\{\sum f_{1}(v_{i}^{\prime})\otimes f_{2}(w_{i}^{\prime}):\sum v_{i}^{\prime}\otimes w_{i}^{\prime}=\sum v_{i}\otimes w_{i}\}.$
It is easy to see that if $f_{1},f_{2}$ are $\supseteq$-morphisms on hyperpairs then $f_{1}\otimes f_{2}$ is a $\supseteq$-morphism. The difficulty with this observation is that for the hyperpairs of greatest interest (for quotient hyperfields) the natural morphisms are $\subseteq$-morphisms, for which $f_{1}\otimes f_{2}$ is not a $\subseteq$-morphism.
@@ 5.3. Tensor extensions of $\preceq$-morphisms and weak morphisms of pre-semialgebra pairs
One can push this further, defining tensor extensions of pre-semialgebras as in Definition 4.22.
@@
Suppose that $\mathcal{A}$ is a weakly admissible $\mathcal{T}$-algebra, $\mathcal{T}^{\prime}$ is a monoid containing $\mathcal{T}$, and $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a pre-semialgebra homomorphism.
@@
    If $\mathcal{T}^{\prime}$ is the disjoint union $\cup_{i\in I}c_{i}\mathcal{T}$ for $c_{i}\in\mathcal{T}^{\prime}$ then $f$ extends naturally to a weak pre-semialgebra morphism (resp.  $\preceq$-morphism) $\tilde{f}:{\mathcal{T}^{\prime}}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{N}_{2}$ given by $\tilde{f}(c_{i}a_{i}\otimes y_{i})=\sum c_{i}\otimes f(a_{i}y_{i}).$
@@
    If $\mathcal{A}_{1}$ is free over $\mathcal{T}$ with base $\{b_{i}:i\in I\},$ then $f$ extends naturally to a weak pre-semialgebra morphism (resp.  $\preceq$-morphism) $\tilde{f}:{\mathcal{T}^{\prime}}\otimes_{\mathcal{T};\operatorname{adm}}\mathcal{M}_{2}\to\mathcal{T}^{\prime}\otimes_{\mathcal{T};\operatorname{adm}}\mathcal{N}_{2}$ given by $\tilde{f}(\sum_{i}a^{\prime}_{i}\otimes b_{i})=\sum_{i}a^{\prime}_{i}\otimes f(b_{i}).$
@@
[…]
$\preceq c_{i}a_{i}\otimes f(y_{i})+c_{i}a_{i}\otimes f(y^{\prime}_{i})=\tilde{f}(c_{i}a_{i}\otimes y_{i})+f(c_{i}a_{i}\otimes y_{i}^{\prime}).$
[…]
@@ 5.4. The adjoint correspondence
The usual proof for modules over rings, say in [22, Proposition 2.10.9] matches values on simple tensors, and then extends the $\subseteq$-morphism $f=f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}$ from its action on simple tensors. This still works for homomorphisms of modules over semirings.
@@
$f_{w}(v_{1}+v_{2})=f\left(\left(\sum v_{i}\right)\otimes w\right)=f\left(\sum v_{i}\otimes w\right)\preceq\sum f\left(v_{i}\otimes w\right),$
and likewise for the other side, so $f\in\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}).$ We get $\Phi\in\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ by declaring $\Phi(f)(w)=f_{w}.$
@@
However, our maps in $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{M}_{3})$ need not be balanced, so might not produce maps of tensor products. In case $\mathcal{M}_{2}$ is free, with base $\mathcal{B}$ over $\mathcal{T},$ we could utilize Proposition 5.10. Define ${\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\mathcal{N})$ to be the multiplicative maps which are 0 except on multiples of base elements.
@@
${\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ is a section of ${\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ when $\mathcal{M}_{2}$ is free over $\mathcal{T}.$
@@
In the other direction, given $g\in{\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3})),$ one defines the map $\psi_{g}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{M}_{3}$ by
@@
Clearly $\Phi(\psi_{g}(w))(v)=0$ unless $w=ab$ for $b\in{\mathcal{B}},$ in which case

### v4

@@ 5.1. Defining $f_{1}\otimes f_{2}$ via simple tensors
Here is the piece that we can salvage in general.
@@
Suppose $\mathcal{M}_{1},\mathcal{N}_{1}$ are $\mathcal{T}$-modules and $\mathcal{M}_{2},\mathcal{N}_{2}$ are left $\mathcal{T}$-modules. If $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ and $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$ are multiplicative maps, then there is a 1-balanced map, which we denote as $f_{1}\times_{\circ}f_{2}:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes\mathcal{N}_{2}$, given by $(b_{1},b_{2})\mapsto f_{1}(b_{1})\otimes f_{2}(b_{2}).$
@@
Given a $(\mathcal{T}_{1},\mathcal{T})$-bimodule prepair $({\mathcal{M}_{1}},{\mathcal{M}_{1}}_{0})$, a $\mathcal{T},\mathcal{T}_{2}$-bimodule prepair $({\mathcal{M}_{2}},{\mathcal{M}_{2}}_{0})$, and a $\mathcal{T}_{1},\mathcal{T}_{2}$-bimodule prepair $(\mathcal{N},\mathcal{N}_{0})$, define $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{N})$ to be the set of $\mathcal{T}_{1},\mathcal{T}_{2}$ multiplicative maps $f:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}$ (as in Remark 5.6) satisfying the property that if $b_{j},b_{i,j}\in\mathcal{M}_{i}$ with $\sum_{j}b_{i,j}\in{\mathcal{M}_{i}}_{0},$ then
@@
Nevertheless, note that $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{N}_{1}\otimes\mathcal{N}_{2})$ need not induce weak morphisms in the tensor product, since there may be some weird sum in $(\mathcal{M}_{1}\times\mathcal{M}_{2})_{0}$.
5.2. The tensor product of morphisms when $\mathcal{M}_{2}$ is free
Proposition 5.5 raises the issue of how to define $(f\otimes g)(v\otimes w)$, hoping to extend it to all of $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$.
@@ 5.2.1. The tensor product of a weak morphism and a homomorphism when $\mathcal{M}_{2}$ is free
When $\mathcal{M}_{2}$ is free as a $\mathcal{T}$-module, we can still define the tensor product of a weak morphism $f_{1}$ and a homomorphism $f_{2}$. Namely, we write tensors uniquely in terms of a base, by taking an arbitrary element $\sum_{i\in I}v_{i}\otimes b_{i}$ as in Corollary 4.16(v), and put $(f_{1}\otimes f_{2})(\sum_{i\in I}v_{i}\otimes b_{i})=f_{1}(v_{i})\otimes f_{2}(b_{i})$. We formulate this as:
Theorem 5.8.
Notation as in Theorem 4.12, suppose $\mathcal{M}_{2}$ is free, with base $\{b_{i}:i\in I\}$. Then for any weak $(\mathcal{T}_{1},\mathcal{T})$-bimodule morphism (resp.  $\preceq$-morphism) $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ and any $(\mathcal{T},\mathcal{T}_{2})$-bimodule homomorphism $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$, there is a weak $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule morphism (resp.  $\preceq$-morphism)
@@
We do it for $\preceq$-morphisms. The proof for weak morphisms is parallel. Define the map $f_{1}\times_{\circ}f_{2}$ as in Proposition 5.5. Then $f_{1}\otimes f_{2}$ is well-defined when we write vectors of $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ uniquely in the form $\sum v_{j}\otimes b_{j}$, and
@@
Corollary 5.9.
We do have functoriality in the residue case under the hypothesis of Lemma 3.12.
Likewise, at times we can extend $\preceq$-morphisms and weak morphisms to tensor extensions.
Theorem 5.10.
Suppose $\mathcal{T}_{1},\mathcal{T}_{2}$ are monoids containing $\mathcal{T}$ where $\mathcal{T}_{1}$ is the disjoint union $\underset{\cdot}{\cup}\,c_{i}\mathcal{T}$, and $\mathcal{M}_{1},\mathcal{M}_{2}$ are $\mathcal{T}$-modules. If $f_{1}:\mathcal{T}_{1}\to\mathcal{T}_{2}$ is a monoid homomorphism and $f_{2}:\mathcal{M}_{1}\to\mathcal{M}_{2}$ is a weak morphism, then there is a weak $\mathcal{T}_{1}$-morphism
$f_{1}\otimes f_{2}:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{1}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{A}_{2}$
given by $(f_{1}\otimes f_{2})(a\otimes b)=f_{1}(a)\otimes f_{2}(b),$ for $a\in\mathcal{T}_{1},\ b\in\mathcal{M}_{1}$
Proof.
$f_{1}\otimes f_{2}$ is well defined because of the hypothesis on $\mathcal{T}_{1},$ so the proof is as in Theorem 4.24(iii). ∎
@@ 5.2.2. The tensor product of two weak morphisms or two $\preceq$-morphisms when $\mathcal{M}_{2}$ is free
When dealing with two weak morphisms or $\preceq$-morphisms, we modify the idea of [20], introducing the empty set and partial maps.
@@ 5.3. $\preceq$-Tensor products of $\preceq$-morphisms
We need to show that $f(\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2})\preceq f(\sum v_{i,1}\otimes w_{i,1})+f(\sum v_{i,2}\otimes w_{i,2})$. But if $\sum v_{i,1}^{\prime}\otimes w_{i,1}^{\prime}=\sum v_{i,1}\otimes w_{i,1}$ and $\sum v_{i,2}^{\prime}\otimes w_{i,2}^{\prime}=\sum v_{i,2}\otimes w_{i,2}$ then
$\sum v_{i,1}^{\prime}\otimes w_{i,1}^{\prime}+\sum v_{i,2}^{\prime}\otimes w_{i,2}^{\prime}=\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2},$
@@
Theorem 5.13 can be applied to hypermodules in view of Remark 3.4, but at cost of introducing the empty set (the $\infty$ element of $\mathcal{P}(\mathcal{H}),$ cf. footnote 3 of Definition 3.1). In other words, $f$ could be undefined on tensors which can be rewritten in inequivalent ways.
@@ 5.3.1. Digression: Hypertensor products of maps
An intriguing (but problematic) way is to introduce a hyperstructure.
@@
Given maps $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$, $i=1,2,$ define $f:=f_{1}\otimes_{\supseteq}f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to{\mathcal{P}}^{\sharp}(\mathcal{N}_{1}\otimes\mathcal{N}_{2})$ by $f(\sum v_{i}\otimes w_{i})=\{\sum f_{1}(v_{i}^{\prime})\otimes f_{2}(w_{i}^{\prime}):\sum v_{i}^{\prime}\otimes w_{i}^{\prime}=\sum v_{i}\otimes w_{i}\}.$
It is easy to see that if $f_{1},f_{2}$ are $\supseteq$-morphisms on hyperpairs then $f_{1}\otimes_{\supseteq}f_{2}$ is a $\supseteq$-morphism. The difficulty with this observation is that for the hyperpairs of greatest interest (for quotient hyperfields) the natural morphisms are $\subseteq$-morphisms.
@@ 5.4. Tensor extensions of $\preceq$-morphisms and weak morphisms of pre-semialgebra pairs
One can push this theory further, taking tensor extensions of pre-semialgebras as in Definition 4.20.
@@
Suppose that $\mathcal{A}$ is a weakly admissible $\mathcal{T}$-semialgebra, $\mathcal{T}^{\prime}$ is a monoid containing $\mathcal{T}$, and $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a weak pre-semialgebra morphism (resp.  $\preceq$-morphism).
@@
    If $\mathcal{A}_{1}$ is free over $\mathcal{T}$ with base $\{b_{i}:i\in I\},$ then $f$ extends naturally to a weak pre-semialgebra morphism (resp.  $\preceq$-morphism) $\tilde{f}:{\mathcal{T}^{\prime}}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{N}_{2}$ given by $\tilde{f}(\sum_{i}a^{\prime}_{i}\otimes b_{i})=\sum_{i}a^{\prime}_{i}\otimes f(b_{i}).$
@@
    If $\mathcal{T}^{\prime}$ is the disjoint union $\underset{\cdot}{\cup}\,c_{i}\mathcal{T}$ for $c_{i}\in\mathcal{T}^{\prime},$ then $f$ extends naturally to a weak pre-semialgebra morphism (resp.  $\preceq$-morphism) $\tilde{f}:{\mathcal{T}^{\prime}}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{N}_{2}$ given by $\tilde{f}(c_{i}a\otimes y)=c_{i}\otimes f(ay).$
@@
$\text{(i)}\quad\tilde{f}\left(\sum_{i}(a_{i}\otimes b_{i})+\sum_{i}(a_{i}^{\prime}\otimes b_{i})\right)=\tilde{f}\bigg(\sum_{i}(a_{i}+a_{i}^{\prime})$ $b_{i}\bigg)$ (5.9)
@@
(ii) Using Theorem 4.24, we have
$\qquad\tilde{f}(c_{i}a\otimes y+$ $c_{i}a^{\prime}\otimes y^{\prime})=\tilde{f}(c_{i}(a\otimes y+a^{\prime}\otimes y^{\prime}))=c_{i}\otimes f(ay+a^{\prime}y^{\prime})$ (5.10)
$\preceq c_{i}a\otimes f(y)+c_{i}a^{\prime}\otimes f(y^{\prime})=\tilde{f}(c_{i}a\otimes y)+\tilde{f}(c_{i}a^{\prime}\otimes y^{\prime}).$
Example 5.16.
Any $\preceq$-morphism (resp. weak morphism) $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ of semiring pairs extends to $\preceq$-morphism (resp. weak morphism) $f:(\mathcal{A}[\lambda],\mathcal{A}_{0}[\lambda])\to(\mathcal{A}^{\prime}[\lambda],\mathcal{A}^{\prime}_{0}[\lambda])$ of the polynomial semiring pairs, taking $\mathcal{T}^{\prime}=\cup_{i\in\mathbb{N}}\mathcal{T}\lambda^{i}.$
@@ 5.5. The adjoint correspondence
The usual proof for modules over rings, say in [22, Proposition 2.10.9] matches values on simple tensors, and then defines the homomorphism $f=f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}$ from its action on simple tensors. This still works for homomorphisms of modules over semirings.
@@
$f_{w}(v_{1}+v_{2})=f\left(\left(\sum v_{i}\right)\otimes w\right)=f\left(\sum v_{i}\otimes w\right)\in{\mathcal{M}_{3}}_{0},$
and likewise for the other side, so $f_{w}\in\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}).$ We get $\Phi\in\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ by declaring $\Phi(f)(w)=f_{w}.$
@@
However, the maps thus obtained in $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{M}_{3})$ need not be balanced, so might not produce maps of tensor products. In case $\mathcal{M}_{2}$ is free, with base $\mathcal{B}$ over $\mathcal{T}_{2},$ we utilize Proposition 5.11. Define ${\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\mathcal{N})$ to be the multiplicative maps which are 0 except on multiples of base elements.
@@
${\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ is a section of ${\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ when $\mathcal{M}_{2}$ is free over $\mathcal{T}_{2}.$
@@
In the other direction, given $g\in{\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3})),$ one defines $\psi_{g}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{M}_{3}$ by
@@
Clearly $\Phi(\psi_{g}(w))(v)=0$ unless $w=ab$ for $a\in\mathcal{T}_{2},$ $b\in{\mathcal{B}},$ in which case